Howard Bowen is a large-scale cotton farmer. The land and machinery he owns has a current market value of $4 million. Bowen owes his local bank $3 million. Last year, Bowen sold $5 million worth of cotton. His variable operating costs were $4.5 million; accounting depreciation was $40,000, although the actual decline in value of Bowen’s machinery was $60,000 last year. Bowen paid himself a salary of $50,000, which is not considered part of his variable operating costs. Interest on his bank loan was $400,000. If Bowen worked for another farmer or a local manufacturer, his annual income would be about $30,000. Bowen can invest any funds that would be derived, if the farm sold, to earn 10 percent annually. (Ignore taxes.)
a. Compute and interpret Bowen’s accounting profits.

b. Compute and interpret Bowen’s economic profits.
1
Expert's answer
2014-06-24T09:19:14-0400
Land and machinery = $4 million owes $3 million
Last year sold $5 million worth of cotton
variable operating costs = $4.5 million;
accounting depreciation = $40,000,
actual decline in value = $60,000
salary = $50,000,
Interest on loan = $400,000
income would be - $30,000.

Bowen can invest any funds that would be derived, if the farm sold, to earn 10 percent annually.

a.Accounting profit consists of revenue minus explicit costs.
Bowen’s accounting profits = Total revenue - Explicit cost = 5,000,000 - (4,500,000 + 40,000 + 50,000 + 400,000) = $10,000

b.Economic profit consists of revenue minus implicit (opportunity) and explicit (monetary) costs.
Bowen’s economic profits = Accounting profit - Implicit cost = 10,000 - (4,000,000*0.1 + 30,000 + (60,000 - 40,000)) = -$440,000